# High Resolution Millimeter Wave Imaging Based on FMCW Radar Systems at W-Band

### v2

@@ Abstract
In this paper, we present a unique 2D high resolution, compact, low-cost, low-weight, and highly accurate millimeter wave imagery system capable of operating in all weather conditions. We describe millimeter wave imaging process in detail and present several novel signal processing methods with their applications. To create the array, we utilize the Synthetic Aperture Radar (SAR) concept. The imagery system presented in this work, can strongly compete with Lidar systems as the resolution limit is at the same level. Furthermore, in contrast to the Lidar systems, our imagery system can operate in heavy rain and dense fog and produce high quality images.
We use our custom-made Frequency Modulated Continuous Wave (FMCW) radar operating at W-band with 33 GHz bandwidth for data collection and present the results.
@@
FMCW radar, high resolution radar imaging, SAR.
@@ I Introduction
Radar imaging is an important topic with many applications that has been used for decades in different areas [1, 2, 3, 4]. The goal in radar imaging is to create 2D or 3D images from the targets located in the radar’s field of view. The image is the representation of the complex reflective coefficient field of the targets. Achieving higher resolution is of high importance in the field of radar imaging. The range resolution is directly related to the bandwidth of the system. Increasing the bandwidth at low frequencies such as L, C, or X bands, that have been used extensively for radar imaging, is considered to be hard to achieve. Moreover, enhancing the number of elements to obtain higher resolution in the azimuth direction, makes the system complex and increases the physical size of the system.
Recently, millimeter, sub-terahertz, and terahertz waves have been used to perform radar imaging [11, 12, 13, 14, 15, 16, 17, 18, 5, 6, 7, 8, 9, 10]. Compared to the systems operating at lower frequencies, the advantage of the systems operating at millimeter wave, sub-terahertz, and terahertz frequencies is that they are light-weight, compact, and have higher bandwidth. Operating with higher bandwidth will allow the system to obtain higher resolution in the range direction. Also, operating at higher frequencies makes it possible to create apertures with large number of elements while they occupy small physical space. This, in turn, yields high resolution in the azimuth direction.
@@ I-A Related Work
The work presented in [11, 12, 13, 14, 15, 16, 17, 18] are based on commercially available FMCW radar [1, 2, 19, 20] chips operating at 77 GHz with 4 GHz bandwidth. High resolution and high quality images have been created. However, due to the fact that the bandwidth of the systems is limited to 4 GHz, therefore, the best range resolution that we can achieve is approximately 5 cm.
In [5], the authors have developed a SiGe-Chip-based 80 GHz Frequency Modulated Continuous Wave (FMCW) radar system with 25.6 GHz bandwidth for high resolution imaging. The work is based on a low power SiGe-MMIC system operating as a FMCW radar. A large aperture has been created synthetically and high resolution images have been generated.
Another interesting work reported in the literature [6, 7] is an imagery system operating at 300 GHz with more than 40 GHz bandwidth. The system has been used in both Synthetic Aperture Radar (SAR) and Inverse SAR (ISAR) modes and has been used to create high resolution images. In [8], a FMCW radar, operating at terahertz frequencies, has been designed to create high resolution images based on Circular SAR (CSAR) technique. In [9], a FMCW radar has been designed at 220 GHz with 8 GHz bandwidth. Then the system has been used to generate high resolution radar images.
The major issue with imagery systems operating at sub-terahertz and terahertz frequencies is low output power and high free-space path-loss which makes them only suitable for near field and very close range applications.
@@ I-B Motivation
In this paper, we present a high resolution imagery system. The system we use in this paper, is our custom-made wide-band radar system. We have chosen the W-band with frequencies ranging from 75 GHz to 108 GHz with 33 GHz bandwidth. Compared to the systems operating at lower frequencies the W-band will allow us to design a compact and light-weight system. In addition, we can reach higher bandwidth which as a result allows to enhance the range resolution limit of the system.
However, increasing the frequency range beyond the W-band will cause issues such as lower output power and higher free-space path-loss which systems presented in [6, 7, 8, 9, 10] suffer from.
Moreover, the system presented in this paper is a custom-made FMCW radar. FMCW radars are well-known for power efficiency and lower complexity compared to the pulsed-radars.
In addition to the system design and image formation, we present several signal processing techniques such as multi-spectral-based imaging, sub-array-based processing, image de-noising, and apodization which distinguishes our work from [11, 12, 13, 14, 16, 17, 18, 5, 6, 7, 8, 9, 10]. The multi-spectral-based imaging subject paves the way to applications such as material characterization. The sub-array-based processing approach provides approximately 3 dB gain which is a remarkable achievement since the origin of the gain is the signal processing unit. We further, describe the system and the imaging process in detail. Moreover, we present image de-noising. We use the state-of-the-art de-noising techniques from the field of image processing and apply it to the reconstructed images. Since the resolution of the reconstructed images is extremely high, therefore, it is plausible to bring sophisticated methods from the filed of image processing in to the field of radar imaging. We, further, utilize apodization technique. The apodization method is performed to reduce the side-lobe levels while preserving the resolution limit of the imagery system [21].
Another major difference between the work we present in this paper and the material reported in [11, 12, 13, 14, 15, 16, 17, 18, 5, 6, 7, 8, 9, 10] is related to the fact that we have collected data from multiple realistic scenarios in different weather conditions. The main purpose is to show the capability of the imagery system presented in this work in producing high resolution images from complex realistic scenarios in different weather conditions such as day, night, rain, and fog.
@@ I-C Organization
The paper has been organized as follows. In Section II, we present the system model along with the imaging algorithm. Section III has been dedicated to the side-lobe reduction using apodization-based methods. In Section IV, we describe the multi-spectral imaging technique. Section V discusses the image de-noising process. In Section VI, we present the sub-array-based processing technique. Finally, section VII has been dedicated to the experimental results followed by the concluding remarks.
@@ II System Model
Fig. 1 illustrates the geometry of the model we are considering in this paper. The array is created synthetically. The imagery system is a single element TX-RX FMCW radar which is moved along the horizontal direction to create a large array based on the SAR concept.
@@
where $f_{c}$ is the carrier frequency and $\sigma_{l}$ is the radar cross section for the $l^{\text{th}}$ point target. The parameter $\beta$ is given as ${b/T}$, where $b$ and $T$ stand for the bandwidth and the chirp time, respectively. Also, $u$ represents the envelope of the signal and is considered to be a rectangular window. Furthermore, $t$ and $\eta$ represent the fast- and slow-time parameters, respectively. Finally, $R_{l}$ is the radial distance between the target and the radar.
Upon taking Fourier transform from (1) with respect to the fast time parameter $t$, we obtain
@@
where $\hat{\sigma}_{l}$ is the estimated reflective coefficient of the $l^{\text{th}}$ point reflector located at the $i^{\text{th}}$ pixel.
@@ III Apodization
The common technique for side-lobe reduction is by applying a window to the data in both range and azimuth directions. However, the side-lobe reduction based on windowing method will reduce the resolution of the system. Apodization is a method to decrease the side-lobe levels without changing the resolution limit of the system. In fact, the resolution of the system remains intact while the side-lobe levels are reduced and will be at the same level of the applied window [21].
Complex dual apodization is a more advanced apodization technique which operates on real and imaginary parts of the signal separately and generates better result compared to the apodization method [21]. In this paper, we will be using both the dual apodization and the complex dual apodization technique to reduce the side-lobe levels while preserving the resolution limit of the imaging system.
@@
In this section, we present multi-spectral-based imaging. The idea is based on partitioning the whole bandwidth into 3 sub-bands. We, then, create 3 separate images and assign red, green, and blue colors to them, respectively. At the end, we display the image in color.
In fact, the sub-bands are created using the IF signal (also known as the beat signal). We divided the beat signal, in fast time domain, into three equal sections and each section represents one-third of the initial signal. Next, each part is used to create an independent image from the scene. Finally, all three images are assigned red, blue, and green colors and the result is plotted.
@@
Fig. 2 shows the process of dividing the entire bandwidth of the system into 3 sub-bands. One of the applications of multi-spectral-based imaging is for materiel characterization in which targets that display the same response over the whole frequency band are presented as white and targets that have frequency-dependent response will be seen as non-white in the processed image.
@@
The image created by the system is corrupted with noise. In this section, we address the image de-noising process. The goal of image de-noising is to remove the noise while prevent image smearing and preserve the fine structures in the image. Image de-noising helps with observing the fine details of the image which have been covered by the noise of the system. Since the imagery system presented in this paper creates incredibly high resolution images, which is comparable to optical images in resolution and quality, therefore, we will be using image de-noising techniques from the filed of image processing.
@@
In (4), $f$ represents the noisy image and $\lambda$ is the hyper-parameter of the optimization problem. Also, $\Omega$ represents the image domain.
@@
To mitigate this issue, we will be solving the Euler-Lagrange equation [24] for the optimization problem given in (4). Moreover, to make the algorithm resilient to edge smearing, we will use an anisotropic version of (4). Consequently, upon using the Euler-Lagrange equation we obtain the following nonlinear elliptic partial differential equation,
@@
In (LABEL:denoising_opt_EL), the parameter K controls the sensitivity to the edges and is a hyper-parameter and * represents the convolution operator. The algorithm presented in (LABEL:denoising_opt_EL) is in fact the nonlinear diffusion method, also known as modified Perona-Malik diffusion technique [25], which is used for image de-noising . We will implement (LABEL:denoising_opt_EL) directly in Python to achieve the de-noised image. One of the main features of the de-noising technique given in (LABEL:denoising_opt_EL) is the ability of the algorithm to preserve the edges and fine details of the image while remove the noise.
@@
In this section, we discuss sub-array-based processing. The goal of the sub-array-based method is to increase the Signal to Noise Ratio (SNR) which is a critical factor for image reconstruction at millimeter wave frequencies. We partition the array into smaller sub-arrays. Then, we perform beam-forming per each sub-array and direct the sub-beams toward a specific spot in the field of view of the imagery system. Next, we utilize all the sub-beams and perform the final beam-forming stage over all the sub-beams that we have created and generate the final image. Fig. 3 illustrates the idea behind the sub-array-based imaging.
Creating sub-arrays this way will increase the SNR which, in turn, will reduce the noise floor and that is the main goal for implementing the sub-array-based approach. The sub-arrays are created with non-zero overlap which forces the noise of the system to become correlated. Therefore, the gain we achieve is not exactly equal to the number of the elements in the sub-arrays. However, we can obtain approximately 3 dB gain in SNR which is a remarkable achievement. It should be noted that, the gain is purely obtained from the signal processing unit and that is what makes this approach valuable.
It is worth mentioning that, increasing the number of elements in the sub-array comes at the cost of narrowing the field of view of the imagery system.
Figure 3: Partitioning the whole array into sub-arrays and direct the sub-beams toward a specific place in the field of view of the imagery system which will be used to perform the final stage of the beam-forming process.
@@
In this section, we present our experimental results. We have designed a FMCW radar at W-band operating at frequencies ranging from 75 GHz to 108 GHz with 33 GHz bandwidth. .
@@
To assure the integrity of the IF signal, a high-pass filter is employed to discard undesirable low-frequency components, with subsequent stages of amplification and filtering enhancing the signal quality. The IF signal is then converted to digital format via an Analog-to-Digital Converter (ADC), making it amenable to perform complex digital processing algorithms.
The purpose of the high pass filter is to remove or decrease the energy of the close-by targets, that are within the minimum range of the radar, as well as blocking the leakage from TX into RX to protect the receiver from saturation and increase the dynamic range of the system.
The ADC block contains the ADC as well as the low pass Anti Aliasing Filter (AAF). The sampling frequency for the ADC is 5 MHz. The chirp time has been set to 5 msec.
@@
Figure 5: a) the impulse response of the system using a corner reflector at 2.05 m radial distance, b) the 3 dB beam-width limit. .
In Fig. 5, we have shown the impulse response of the system using a corner reflector located at 2.05 m radial distance.
To perform the image reconstruction process, we have chosen frequencies ranging from 78 GHz to 102 GHz which is the flat part of the spectrum that the system provides and is equivalent to 24 GHz bandwidth. The 24 GHz bandwidth will provide us with 6.3 mm resolution in the range direction. The length of the array, for all the experiments that we present in this section, is 40 cm.
Fig. 6 illustrates the power spectrum of the system over the entire bandwidth ranging from 75 GHz to 108 GHz. It is clear from Fig. 6 that we have large power fluctuation in the beginning and at the end of the spectrum. For the imaging process, we have chosen the part corresponding to range of frequencies from 78 GHz to 102 GHz which is equivalent to 24 GHz bandwidth.
@@
Figure 7: The radar system as well as the motorized scanner to create the synthetic aperture.
Fig. 7 illustrates the radar system as well as the motorized scanner to create the synthetic aperture.

### v3

@@ Abstract
In this paper, we present a unique 2D high resolution, compact, low-cost, low-weight, and highly accurate millimeter wave imagery system capable of operating in all weather conditions. We describe millimeter wave imaging process in detail and present several novel signal processing methods with their applications. To create the array, we utilize the Synthetic Aperture Radar (SAR) concept. The imagery system presented in this paper, can strongly compete with Lidar systems as the resolution limit is at the same level. Furthermore, in contrast to the Lidar systems, our imagery system can operate in heavy rain and dense fog and produce high quality images.
Finally, we utilize our wide-band custom-made Frequency Modulated Continuous Wave (FMCW) radar, which operates at W-band with 33 GHz bandwidth, for data collection and present the results.
@@
FMCW radar, high resolution imaging, millimeter wave imaging, SAR.
@@ I Introduction
Radar imaging is an important topic with numerous applications which has been used in different areas over the past decades[1, 2, 3, 4]. The goal in radar imaging is to create 2D or 3D images from the targets located in the radar’s field of view. The image is the representation of the complex radar cross section of the targets. In the field of radar imaging, creating high resolution images is of utmost importance. The range resolution is directly related to the bandwidth of the system. However, increasing the bandwidth of the system at low frequencies such as L, C, or X bands, that have been utilized extensively for radar imaging, is challenging. Moreover, enhancing the number of elements to obtain higher resolution in the azimuth direction, enhances the complexity of the system and increases its physical size as well.
Recently, millimeter, sub-terahertz, and terahertz waves have been used to perform radar imaging [11, 12, 13, 14, 15, 16, 17, 5, 6, 7, 8, 9, 10]. Compared to the systems that operate at lower frequencies, the advantage of the systems which operate at millimeter wave, sub-terahertz, and terahertz frequencies is that they are light-weight, compact, and have higher bandwidth. Operating with higher bandwidth will allow the system to obtain higher resolution in the range direction. In addition, operating at higher frequencies makes it possible to create apertures with large number of elements while the array still occupies a small physical space. This, in turn, yields higher resolution in the azimuth direction.
@@ I-A Related Work
The work presented in [11, 12, 13, 14, 15, 16, 17] are based on commercially available Frequency Modulated Continuous Wave (FMCW) radar chips that operate at 77 GHz with 4 GHz bandwidth [1, 2, 18, 19]. High resolution and high quality images have been created. However, due to the fact that the bandwidth of the systems is limited to 4 GHz, therefore, the best range resolution that we can achieve is approximately 5 cm.
In [5], the authors have developed a SiGe-Chip-based 80 GHz FMCW radar system with 25.6 GHz bandwidth for high resolution imaging. The work is based on a low power SiGe-MMIC system operating as a FMCW radar. A large aperture has been created synthetically and high resolution images have been generated.
Another interesting work reported in the literature [6, 7] is an imagery system operating at 300 GHz with more than 40 GHz bandwidth. The system has been used in both Synthetic Aperture Radar (SAR) and Inverse SAR (ISAR) modes and has been utilized to create high resolution images. In [8], a FMCW radar, operating at terahertz frequencies, has been designed to create high resolution images based on Circular SAR (CSAR) technique. In [9], a FMCW radar has been designed at 220 GHz with 8 GHz bandwidth. Then, the system has been used to generate high resolution radar images.
The major issue with imagery systems operating at sub-terahertz and terahertz frequencies is low output power and high free-space path-loss which makes them only suitable for near-field and close-range applications.
@@ I-B Motivation
In this paper, we present a high resolution imagery system. The system which we present in this paper, is our custom-made wide-band FMCW radar system. We have selected the W-band with frequencies ranging from 75 GHz to 108 GHz with 33 GHz bandwidth. Compared to the systems operating at lower frequencies, the W-band will make it possible to design a compact and light-weight system. In addition, we can reach higher bandwidth which as a result allows to enhance the range resolution limit of the system.
However, increasing the frequency range beyond the W-band will cause issues such as lower output power and higher free-space path-loss which systems presented in [6, 7, 8, 9, 10] suffer from. Moreover, the system presented in this paper is a custom-made FMCW radar. FMCW radars are well-known for power efficiency and lower complexity compared to the pulsed-radars.
In addition to the system design and image formation, we present several signal processing techniques such as multi-spectral-based imaging, sub-array-based processing, image de-noising, and apodization which distinguishes our work from [11, 12, 13, 14, 15, 16, 17, 5, 6, 7, 8, 9, 10]. The multi-spectral-based imaging subject paves the way to applications such as material characterization. The sub-array-based processing approach provides approximately 3 dB gain which is a remarkable achievement since the origin of the gain is the signal processing unit. We, further, describe the system and the imaging process in detail. Moreover, we present image de-noising. We use the state-of-the-art de-noising techniques from the field of image processing and apply it to the reconstructed images. Since the resolution of the reconstructed images is extremely high, therefore, it is plausible to bring sophisticated methods from the field of image processing in to the field of radar imaging. In addition, we utilize apodization technique. The apodization method is performed to reduce the side-lobe levels while preserving the resolution limit of the imagery system [20].
Another major difference between the work we present in this paper and the material reported in [11, 12, 13, 14, 15, 16, 17, 5, 6, 7, 8, 9, 10] is related to the fact that we have collected data from multiple realistic scenarios in different weather conditions. The main purpose is to show the capability of the imagery system, presented in this work, in producing high resolution images from complex realistic scenarios in different weather conditions such as day, night, rain, and fog.
@@ I-C Organization
The paper has been organized as follows. In Section II, we present the system model along with the imaging algorithm. Section III has been dedicated to the side-lobe reduction using apodization-based methods. In Section IV, we describe the multi-spectral-based imaging technique. Section V discusses the image de-noising process. In Section VI, we present the sub-array-based processing method. Finally, section VII has been dedicated to the experimental results which is followed by the concluding remarks.
@@ II System Model
Fig. 1 illustrates the geometry of the model that we study in this paper. The array is created synthetically. The imagery system is a single element TX-RX FMCW radar which is moved along the horizontal direction to create a large array based on the SAR concept.
@@
where $f_{c}$ is the carrier frequency and $\sigma_{l}$ is the radar cross section for the $l^{\text{th}}$ point target. The parameter $\beta$ is given as ${b/T}$, where $b$ and $T$ stand for the bandwidth and the chirp-time, respectively. Also, $u$ represents the envelope of the signal and is considered to be a rectangular window. Furthermore, $t$ and $\eta$ represent the fast- and slow-time parameters, respectively. Finally, $R_{l}$ is the radial distance between the target and the radar.
Upon taking Fourier transform with respect to the fast-time parameter $t$ from (1), we obtain
@@
where $\hat{\sigma}_{l}$ is the estimated complex valued radar cross section of the $l^{\text{th}}$ point reflector located at the $i^{\text{th}}$ pixel.
@@ III Apodization
The common technique for side-lobe reduction in radar imaging, is by applying a window to the data in both the range and azimuth directions. However, the side-lobe reduction based on windowing method will reduce the resolution of the system. Apodization is a method which is utilized to decrease the side-lobe levels without changing the resolution limit of the system. In fact, the resolution of the system remains intact while the side-lobe levels are reduced and will be at the same level of the applied window [20].
Complex dual apodization is yet a more advanced apodization technique which operates on real and imaginary parts of the signal separately and generates better results compared to the apodization method [20]. In this paper, we will be using both the dual apodization and the complex dual apodization techniques to reduce the side-lobe levels while preserving the resolution limit of the imagery system.
@@
In this section, we present multi-spectral-based imaging. The concept is based on partitioning the entire bandwidth into 3 sub-bands. We, then, create 3 separate images and assign red, green, and blue colors to them, respectively. At the end, we display the image in color.
In fact, the sub-bands are created using the IF signal (also known as the beat signal). We divide the beat signal, in fast-time domain, into three equal sections where each section represents one-third of the initial signal. Next, each part is used to create an independent image from the scene. Finally, all three images are assigned red, blue, and green colors, respectively, and the result is plotted.
@@
Fig. 2 shows the process of dividing the entire bandwidth of the system into 3 sub-bands. One of the applications of multi-spectral-based imaging is for materiel characterization in which targets that display the same response over the entire frequency band are presented as white and targets that have frequency-dependent response will be seen as non-white in the processed image.
@@
The image created by the system is contaminated by the noise of the system. In this section, we address the image de-noising process. The goal of image de-noising is to remove the noise while the edges as well as the fine structures in the image are preserved. Furthermore, image smearing and blurring will not occur. Image de-noising assists with observing the fine details of the image which have been covered by the noise of the system. Since the imagery system, presented in this paper, creates incredibly high resolution images, which is comparable to optical images in resolution and quality, therefore, we will be using image de-noising techniques from the filed of image processing.
@@
In (4), $f$ represents the noisy image and $\lambda$ is the hyper-parameter of the optimization problem. Furthermore, $\Omega$ represents the image domain.
@@
To mitigate this issue, we will be solving the Euler-Lagrange equation [23] for the optimization problem given in (4). Moreover, to make the algorithm resilient to edge smearing, we will utilize an anisotropic version of (4). Consequently, upon using the Euler-Lagrange equation we obtain the following nonlinear elliptic partial differential equation,
@@
In (LABEL:denoising_opt_EL), the parameter K controls the sensitivity to the edges and is a hyper-parameter and ()(*)( ∗ ) represents the convolution operator. The algorithm presented in (LABEL:denoising_opt_EL) is, in fact, the nonlinear diffusion method, also known as modified Perona-Malik diffusion technique [24], which is used for image de-noising.
We will implement (LABEL:denoising_opt_EL) directly in Python to mitigate the effect of the noise. One of the main features of the de-noising technique given in (LABEL:denoising_opt_EL) is the ability of the algorithm to preserve the edges and fine details of the image while being able to remove the effect of the noise.
@@
In this section, we discuss sub-array-based processing. The goal of the sub-array-based method is to increase the Signal to Noise Ratio (SNR) which is a crucial factor for image reconstruction at millimeter wave frequencies. To begin with, we partition the array into smaller sub-arrays. Then, we perform beam-forming per each sub-array and direct the sub-beams toward a specific spot in the field of view of the imagery system. Next, we utilize all the sub-beams and perform the final beam-forming stage over all the sub-beams that we have created and generate the final image. Fig. 3 illustrates the sub-array-based imaging concept.
The process of creating the sub-arrays will increase the SNR which, in turn, will reduce the noise floor which is, indeed, the main goal for the implementation of the sub-array-based approach. Nevertheless, since the sub-arrays are created with non-zero overlap, which forces the samples of the noise to become correlated, therefore, the gain that we achieve by performing the sub-array-based processing, is not exactly equal to the number of the elements in the sub-arrays. However, we can still obtain approximately 3 dB gain in SNR which is considered to be a remarkable achievement. It should be noted that, the gain is purely obtained from the signal processing unit and that is what makes this approach valuable. In other words, this approach does not increase the complexity of the imagery system.
It is worth mentioning that, increasing the number of elements in the sub-array results in narrower field of view for the imagery system.
Figure 3: Partitioning the entire array into sub-arrays and directing the sub-beams toward a specific place in the field of view of the imagery system which will be used to perform the final stage of the beam-forming process.
@@
In this section, we present the experimental results. We have designed a FMCW radar at W-band which operates at frequencies ranging from 75 GHz to 108 GHz with 33 GHz bandwidth.
@@
To assure the integrity of the IF signal, a high-pass filter is employed to discard undesirable low-frequency components, with subsequent stages of amplification and filtering enhancing the signal quality. The IF signal is then converted to digital format via an Analog-to-Digital Converter (ADC), making it amenable to perform complex digital processing algorithms. The purpose of the high pass filter is to remove or decrease the energy of the close-by targets that are within the minimum range of the radar. Furthermore, the high pass filter blocks the leakage from the TX into RX and this way it protects the receiver from saturation and it also increases the dynamic range of the system. The ADC block contains the ADC as well as the low pass Anti Aliasing Filter (AAF). The sampling frequency for the ADC is 5 MHz. The chirp time has been set to 5 msec.
@@
Figure 5: a) the impulse response of the system using a corner reflector at 2.05 m radial distance, b) the 3 dB beam-width limit.
In Fig. 5, we have shown the impulse response of the system using a corner reflector located at 2.05 m radial distance. To perform the image reconstruction process, we have selected the frequency components that rang from 78 GHz to 102 GHz which is the flat part of the spectrum that the system provides and is equivalent to 24 GHz bandwidth. The 24 GHz bandwidth will provide us with 6.3 mm resolution in the range direction. The length of the array, for all the experiments that we present in this section, is 40 cm.
Fig. 6 illustrates the power spectrum of the system over the entire bandwidth ranging from 75 GHz to 108 GHz. From Fig. 6, it is clear that the system is experiencing large power fluctuation in the beginning and at the end of the spectrum. For the imaging process, we have selected the part of the spectrum corresponding to the range of frequencies from 78 GHz to 102 GHz which is equivalent to 24 GHz bandwidth.
@@
Figure 7: The radar system along with the motorized scanner to create the synthetic aperture.
Fig. 7 illustrates the radar system along with the motorized scanner to create the synthetic aperture.